### v2

# StrucText-Eval: An Autogenerated Benchmark for Evaluating Large Language Model’s Ability in Structure-Rich Text Understanding

Abstract
Given the substantial volumes of structured data held by many companies, enabling Large Language Models (LLMs) to directly understand structured text in non-structured forms could significantly enhance their capabilities across various business scenarios. To this end, we propose evaluation data generation method for assessing LLM’s ability in understanding the structure-rich text, which generates structured data of controllable complexity based on manually crafted question templates and generation rules. Building on this generation method, we introduce StrucText-Eval, a benchmark comprising 6,032 questions across 8 different structured languages and 29 specific tasks. Furthermore, considering human proficiency in rule-based tasks, we also present StrucText-Eval-Hard, which includes 3,016 questions designed to further examine the gap between LLMs and human performance. Results indicate that the best-performing LLM currently achieve an accuracy of 65.0% on StrucText-Eval-Hard, while human accuracy reaches up to 95.7%. Moreover, while fine-tuning using StrucText-Eval can enhance existing LLMs’ understanding of all structured languages, it does not necessarily improve performance across all task types. The benchmark and generation codes are open sourced in https://github.com/MikeGu721/StrucText-Eval
1 Introduction
Structured data, often represented by various structured languages such as JSON [19], YAML [8], ORG [11], or Markdown [13], Latex [15] etc., has consistently been central to corporate data strategies due to its ability to capture, store, and analyze essential information systematically. The inherent benefits of structured data lie in its standardized format and high degree of organization, which facilitates efficient data querying and machine processing, clearly surpassing the inherent chaos of unstructured data. However, with the advancement of large language models (LLMs) [1, 26, 27, 22], there has been a significant shift towards the effective utilization of unstructured data, attributed to the LLMs’ capacity to comprehend and generate complex and nuanced semantics within such data [3]. Considering that structured data can be directly presented in an unstructured format, it makes us wonder: whether it is possible to rely on LLMs to interpret structured data through unstructured format directly.
Current LLMs researches have addressed their comprehension of structure-rich text, with a primary focus on the understanding of limited categories: Graphs [9, 18, 14, 24, 6], Tables [21, 4, 17] and JSON [5, 23]. However, these categories do not encompass all potential use cases of structure-rich text. For instance, scenarios requiring direct understanding of articles in Latex or Markdown formats, data in YAML or ORG formats, or various custom structured languages are not adequately covered. Moreover, existing benchmarks often rely on manually annotated data for evaluation, which limits the development of robust evaluation frameworks and potentially facilitates model cheating [31].
In response, we propose a method for automatically generating evaluation data to assess models’ capabilities in understanding structure-rich text. This method is applied to 8 structured languages, as shown in Fig. 1, across 29 specific tasks, enabling data generation with controllable difficulty by adjusting the depth of structured nesting. Based on this method, we further introduce the Structure-Rich Text Evaluation Benchmark (StrucText-Eval), a comprehensive benchmark with 6,032 samples designed to evaluate the proficiency of LLMs in deciphering embedded structures within input text. StrucText-Eval aims to determine whether LLMs understand raw structural tags, execute logical inferences based on the decoded semantics of these symbols, and construct their responses according to instruction requirements. Considering that humans excel at understanding structured expressions, we further propose StrucText-Eval-Hard, a subset of StrucText-Eval consisting of 3,012 samples with the longest questions, highlighting the disparity between LLMs and human understanding of structured data.
The experiments show that StrucText-Eval significantly challenges current LLMs. The highest accuracy achieved by LLMs using various prompt methods was 78.4%, dropping to 65.0% on StrucText-Eval-Hard. In contrast, human participants reached 95.7% accuracy on StrucText-Eval-Hard, underscoring the inadequacy of existing LLMs in processing complex structural information. Utilizing StrucText-Eval data enhances the capabilities of current closed-source LLMs and improves open-source LLMs by fine-tuning them with this benchmark. This finding confirms StrucText-Eval’s critical role in evaluating and optimizing model performance. However, more than using the training data from StrucText-Eval alone is required to surpass GPT4-Turbo’s performance on this test suite. Additionally, not all tasks can be improved solely by training on the relevant dataset. Therefore, StrucText-Eval provides a comprehensive evaluation framework for assessing LLMs’ ability to understand structured text, leaving room for further improvement.
We present the following contributions to our research: (1) We developed a method for generating evaluation data for eight structured languages. This method allows control over the complexity of the generated evaluation data by manipulating the quantity of structural nesting. (2) We introduce StrucText-Eval, a comprehensive benchmark comprising 6,032 samples across eight structured languages and 29 evaluation tasks. This benchmark is designed to assess LLMs’ capabilities in understanding structured data comprehensively. (3) We have conducted extensive experiments, including evaluations of existing large models’ capabilities in understanding structure-rich text and enhancements to the capacity of open-source LLMs in this area. Our findings indicate a significant gap between current LLMs and humans in understanding structure-rich text.
2 Related Work
2.1 Structural Text Understanding Enhanced Models
Previous explorations over structural information focused on tasks related to programming languages and graphs. Mou et al. [16] built a tree-based convolutional neural network that outperformed language models on functionality classification tasks and pattern detection tasks over programming languages. Wang et al. [29] outperformed state of the art (SOTA) on code clone detection benchmarks, applying graph neural networks on abstract syntax trees (ASTs) representation augmented by control and data flow edges, which is different in architecture from language models curated for natural language tasks. Zhou et al. [32] built a graph neural network over rich node representations of code and gained SOTA on vulnerability identification in 4 large-scaled open-source C projects. Perozzi et al. [18] proposed a dedicated graph neural network (GNN) to encode graph into embedding space of LLMs, and used LLMs as an decoder to solve the graph tasks in GraphQA [9]. Chen et al. [6] explored LLMs’ potential in learning graphs, focusing on node classification tasks, using LLMs as enhancers for GNN predictor and directly as predictor. Chen et al. [5] disclosed the employment of formats other than natural language in CoT reasoning and multi-agent communication significantly enhances model performance while also decreasing token usage on reasoning tasks.
2.2 Structural Text Understanding Evaluation Datasets
The similar datasets dedicating to structure information understanding are scarce. GraphQA [9] is a benchmark for assessing LLMs on reasoning over graph data, composed of 7 simple graph tasks, with exploration on methods in encoding graph-structured data as text. Struc-Bench [25] collect data from open source to evaluate if LLM can generate complex structured tabular text. Sui et al. [21] consists of 8 tasks on table-formatted input texts, i.e., input data organized in row-column structures, with studying in input design and prompt enhancement. Pasupat and Liang [17] exposed a dataset of 22,033 complex questions on Wikipedia tables and proposed a parsing algorithm for answering complex questions on semi-structured tables. For tabular data, Shwartz-Ziv and Armon [20] showed that deep models were not a good solution; Gorishniy et al. [12] designed two specialized baseline architecture for tabular data and concluded that universally superior solution does not exist. Wang et al. [28] is another graph-based problem solving benchmark, with graph structure described with natural language. Fatemi et al. [9] and Sui et al. [21] include comparison of encoding methods for structure information, i.e., graphs and tables, respectively, showing that LLMs performance varies on different encodings and the relatively vacancy on LLM’s understanding of structure-rich texts input.
Our dataset diverges from previous work in two significant ways. Firstly, our tasks exclusively concentrate on analyzing structural information and executing dependent inference, unlike prior approaches [16, 32] that employ structural data primarily as an auxiliary tool for classification or semantic comprehension. StrucText-Eval focus on symbolic structure by erasing content semantics from input texts, preserving only structural details, which renders the texts semantically obscure and the tasks semantically agnostic. Secondly, our benchmark extends beyond traditional graph reasoning and table information retrieval to include a broader spectrum of structure-rich text types. It also incorporates comprehensive structure-based inference tasks that are tailored to specific input formats.
3 StrucText-Eval Construction
3.1 Structure-Rich Texts Taxonomy
[…]
3.2 Dataset Generation
An example of JSON’s PathCompose is shown in Fig. 2 to illustrate the dataset generation process. The generation process mainly entails constructing an abstract structure tree, manually drafting question templates, and developing corresponding answer discovery algorithms. The primary objective of constructing the abstract structure tree is establishing a hierarchically nested structured data template. The complexity of the sample increases when more nodes are set in this tree.
Eight task categories have been delineated for eight languages, as detailed in Fig. 0(a). 29 rules and question templates have been formulated for these tasks, with the specific rule templates detailed in Sec. F in the Appendix. Each sample in the dataset comprises four main fields: “Reference”, “Question”, “Requirement” and “Answer”. We give examples for each language and task in Sec. E in the Appendix.
Task Name | Abbr. | Task Description |
Syntax | SY | Focuses on detecting structural errors in data formats such as JSON, XML, and YAML. |
PathWalk | PW | Focuses on extracting specific sections or subsections from structured documents such as org, LaTeX, or markdown files. |
TextRetrieval | TR | Assesses the ability to extract specific information from various document formats, including text content and image filenames. |
Statistic | ST | Concentrates on statistical queries to calculate the number of employees meeting specific salary conditions. |
Join | JO | Assesses the ability to filter data sets that meet specific criteria by combining multiple tables in a database through SQL queries. |
Tree.Height | TH | Evaluates calculating the height of the longest path from the root node to any leaf node in a tree structure. |
Node.Depth | ND | Assesses the depth of any node in a tree structure relative to the root node. |
PathCompose | PC | Evaluates reasoning of paths and multi-level data indexing within hierarchical or tree-like structures. |
3.3 Statistic Information
StrucText-Eval has assembled a dataset comprising 6,032 samples, with each of the 29 specific tasks for eight languages containing 208 samples. These tasks are divided into “simple” and “hard” categories based on data length and complexity, as depicted in Fig. 3(a). The dataset is partitioned into training and testing sets at a 9:2 ratio, which supports extensive training while also enabling the evaluation of model generalization on significant volumes of unseen data. Detailed statistics regarding the number of samples, lengths, and complexity levels across all tasks, languages and difficulties are detailed in Tab. 3 in Sec. C in Appendix.
4 Experiment Setup
To evaluate LLMs’ current capability of processing structure-rich text and executing dependent inference, we conducted a series of experiments using StrucText-Eval in various settings. Our study utilizes both prompt-based and finetuning methods to analyze the performance variations.
4.1 Models
We tested five Close-Source LLMs accessed through public APIs: GPT4, GPT4 Turbo, MINIMAX, SPARK, and ERNIE. Besides these API-Based models, we also finetuned Llama-7B-v1 under several conditions with partially frozen parameters. The API calling and finetuning setting is listed in Appendix C.
4.2 Prompt-based Method
We assessed the influence of prompt design on LLM performance by employing four distinct prompt settings, and the detailed implementation of these prompts can be found in Tab. LABEL:tab:prompt in the Appendix. Naive: This setting involves appending the question directly after the input text, serving as a straightforward task formulation. CoT: In this configuration, the LLM generates its thinking process before directly response to the question, aimed at facilitating task resolution [30]. Few-Shot Demonstration: This approach leverages the principles of few-shot learning [10], with variations including one and two example instances to guide the model. Rule Hint: This setup incorporates manually crafted, step-by-step guidance appended after the input text to aid the inference process.
4.3 Finetune-based Method
We further explored the efficacy of finetuning OpenLLaMA 7B v1 using various subsets of data. Our finetuning configurations are described as follows: Base: Utilizes the core StrucText-Eval dataset, referred to as the base StrucText-Eval setting. 2-Shot: Implements a few-shot (2-shot) learning strategy within the StrucText-Eval framework. Rule Hint: Adapts StrucText-Eval by appending a new task-specific rule hint to each sample, effectively redefining the input text as a combination of the question, answer, and original input, with the answer delineated as a step-by-step procedure. Mix: is a variant of Base with half of samples replaced by samples from ShareGPT111https://huggingface.co/datasets/shibing624/sharegpt_gpt4.
4.4 Evaluation Method
This paper employs three distinct evaluation methodologies: GPT-Judge and RougeL. GPT-Judge utilizes GPT4-turbo to assess whether the LLM output semantically aligns with the ground truth. RougeL measures the degree of character similarity between them. Given that the RougeL expects LLM to output no superfluous content, it imposes additional demands on the LLM’s ability to adhere to instructions, potentially distorting the evaluation results. Therefore, we primarily present the results of GPT-Judge in the main text, while the outcomes of RougeL are listed in Sec. G in the Appendix.
5 Analysis
Model | Prompt | Languages | Tasks | all | ||||||||||||||
[…]
Ernie-turbo | Base | 27.4 | 0.2 | 0.0 | 0.0 | 13.0 | 2.6 | 14.3 | 24.4 | 2.8 | 22.3 | 23.2 | 11.3 | 0.5 | 11.1 | 5.4 | 1.8 | 12.5 |
CoT | 19.4 | 0.1 | 0.1 | 0.2 | 10.9 | 1.5 | 13.4 | 19.7 | 0.8 | 6.7 | 26.6 | 6.2 | 1.4 | 6.2 | 1.8 | 0.0 | 7.2 | |
1-shot | 25.0 | 0.0 | 4.2 | 8.7 | 21.9 | 12.5 | 8.3 | 25.0 | 12.5 | 20.0 | 8.3 | 19.4 | 0.0 | 0.0 | 25.0 | 0.0 | 15.2 | |
2-shot | 12.5 | 12.5 | 12.5 | 8.3 | 21.9 | 20.8 | 16.7 | 5.0 | 4.2 | 10.0 | 4.2 | 18.3 | 25.0 | 0.0 | 25.0 | 25.0 | 13.4 | |
Rule Hint | 32.5 | 0.0 | 0.0 | 0.0 | 9.4 | 0.0 | 12.5 | 32.5 | 8.3 | 20.0 | 25.0 | 14.4 | 0.0 | 6.2 | 0.0 | 0.0 | 13.8 | |
[…]
CoT | 82.5 | 58.3 | 43.8 | 31.3 | 17.3 | 16.6 | 60.4 | 62.5 | 29.2 | 55.4 | 81.6 | 67.3 | 17.1 | 24.2 | 12.8 | 0.0 | 39.1 | |
1-shot | 80.0 | 50.0 | 58.3 | 50.0 | 37.5 | 50.0 | 62.5 | 85.0 | 29.2 | 67.5 | 91.7 | 68.3 | 25.0 | 18.8 | 37.5 | 98.0 | 61.6 | |
2-shot | 90.0 | 58.3 | 54.2 | 54.2 | 50.0 | 45.8 | 62.5 | 82.5 | 41.7 | 70.0 | 91.7 | 70.2 | 50.0 | 31.2 | 12.5 | 96.7 | 65.1 | |
Rule Hint | 77.5 | 37.5 | 54.2 | 50.0 | 40.6 | 16.7 | 66.7 | 62.5 | 8.3 | 55.0 | 91.7 | 65.4 | 12.5 | 25.0 | 50.0 | 0.0 | 53.0 | |
[…]
CoT | 52.5 | 20.8 | 4.3 | 16.7 | 6.2 | 24.9 | 65.8 | 66.4 | 41.7 | 45.3 | 83.3 | 36.0 | 0.0 | 12.6 | 0.0 | 54.1 | 39.8 | |
1-shot | 27.5 | 8.3 | 4.2 | 4.2 | 3.1 | 4.2 | 33.3 | 37.5 | 0.0 | 12.5 | 62.5 | 18.3 | 0.0 | 0.0 | 0.0 | 12.5 | 17.2 | |
2-shot | 30.0 | 25.0 | 41.7 | 12.5 | 0.0 | 66.7 | 66.7 | 62.5 | 37.5 | 22.5 | 91.7 | 36.5 | 0.0 | 0.0 | 50.0 | 75.0 | 37.9 | |
Rule Hint | 42.5 | 4.2 | 8.3 | 0.0 | 0.0 | 8.3 | 41.7 | 7.5 | 8.3 | 5.0 | 70.8 | 13.5 | 0.0 | 0.0 | 0.0 | 0.0 | 15.1 | |
[…]
CoT | 98.4 | 66.7 | 67.8 | 56.5 | 71.9 | 37.5 | 87.5 | 87.5 | 75.0 | 84.6 | 87.5 | 76.9 | 62.5 | 68.8 | 37.1 | 25.8 | 74.9 | |
1-shot | 95.0 | 66.7 | 70.8 | 78.3 | 68.8 | 54.2 | 79.2 | 97.5 | 87.5 | 80.0 | 91.7 | 84.0 | 75.0 | 50.0 | 37.5 | 50.0 | 78.9 | |
2-shot | 95.0 | 79.2 | 65.2 | 75.0 | 87.5 | 54.2 | 87.5 | 97.5 | 79.2 | 82.1 | 95.8 | 84.6 | 96.0 | 75.0 | 50.0 | 62.5 | 82.7 | |
Rule Hint | 97.5 | 66.7 | 65.2 | 50.0 | 65.6 | 45.8 | 83.3 | 95.0 | 83.3 | 80.0 | 91.7 | 75.7 | 50.0 | 56.2 | 75.0 | 12.5 | 74.5 | |
[…]
CoT | 81.6 | 57.4 | 52.0 | 45.2 | 63.6 | 43.8 | 78.8 | 79.2 | 61.1 | 63.8 | 76.2 | 62.8 | 52.7 | 74.6 | 58.3 | 7.1 | 62.6 | |
1-shot | 92.5 | 66.7 | 62.5 | 62.5 | 59.4 | 58.3 | 70.8 | 82.5 | 66.7 | 77.5 | 70.8 | 78.8 | 37.5 | 50.0 | 75.0 | 37.5 | 71.6 | |
2-shot | 87.5 | 62.5 | 66.7 | 75.0 | 68.8 | 54.2 | 75.0 | 92.5 | 83.3 | 75.0 | 75.0 | 80.8 | 50.0 | 68.8 | 50.0 | 37.5 | 75.0 | |
Rule Hint | 82.5 | 62.5 | 62.5 | 50.0 | 78.1 | 29.2 | 70.8 | 87.5 | 66.7 | 60.0 | 70.8 | 76.9 | 25.0 | 93.8 | 50.0 | 12.5 | 68.5 |
Model | Prompt | ||||
Base | CoT | 1-shot | 2-shot | Rule Hint | |
Ernie-turbo | 4.9 | 2.4 | 6.7 | 7.9 | 5.5 |
Minimax | 31.3 | 20.7 | 40.3 | 44.3 | 35.7 |
Spark | 16.3 | 22.4 | 10.7 | 25.3 | 11.5 |
GPT4 | 65.0 | 59.3 | 59.9 | 64.1 | 57.2 |
GPT4-turbo | 54.2 | 50.7 | 49.4 | 53.8 | 49.4 |
Human | 95.7 | - | - | - | - |
5.1 Close Sourced LLMs’ Performance
In this section, we analyze the performance of closed-source LLMs in various tasks involving structure-rich text and multiple languages. Table 5 displays the overall performance on the complete StrucText-Eval dataset, while Table 2 compares both LLM and human performance specifically on the StrucText-Eval-Hard subset.
5.1.1 GPT4 Series Outperform All Other LLMs in Structure-Rich Text Understanding
The experiments revealed that GPT4 consistently outperformed other models across most tasks and languages. Specifically, GPT-4 achieved an average accuracy of 78.4 in all languages and tasks, significantly higher than the next best model, GPT4-turbo, which had an accuracy of 69.3. The marked disparity in performance between GPT4 Series and other LLMs underscores its robustness in tasks requiring deep semantic understanding and complex data manipulation. This demonstrates the potential of advanced models like GPT-4 in environments demanding high accuracy and versatility while also indicating the need for further development in other models to close the performance gap.
5.1.2 Current LLMs Cannot Truly Understand Structure-Rich Expressions
We conclude that although the performance of LLMs improves with more training data, they still fall short of truly comprehending structured language expressions, as evidenced by two observations. Firstly, evaluation results indicate that models from various companies exhibit distinct performance rankings across different languages. Notably, there is consistently lower performance in Tree structures, which are entirely composed of custom languages. This suggests that current LLMs are incapable of understanding complex, custom-structured expressions. Furthermore, results from utilizing Rule Hints generally underperform relative to the Base Prompt, implying that providing models with additional parsing rules may impair performance. In contrast, the 2-shot method consistently enhances the overall performance of LLMs, indicating that while LLMs are able to grasp individual examples, they do not fully understand the underlying structures or the rationale behind them.
5.1.3 Procedural Task is Watershed For LLMs
An in-depth analysis of the performance data from several LLMs demonstrates a clear demarcation in their ability to handle procedural tasks, which require navigating and manipulating structured data. Tasks emphasising procedural complexity, such as “PathCompose” and “Tree.Height”, present significant challenges to these models. For example, in the “PathCompose” task, GPT-4 shows a superior performance with an accuracy rate of 79.2% in its base setting, compared to Minimax, which achieves only 25.0%. Similarly, in the “Tree.Height” task, GPT-4 maintains a robust performance with 37.5% accuracy, whereas Minimax struggles, recording 0% in its base prompt setting.
5.1.4 Impact of Data Length and Complexity on LLM Performance
Fig. 5 illustrates how variations in Data Length and Complexity affect the performance of GPT4-Turbot. Specifically, we quantify Data Length and Complexity using the character length of References and the number of Nodes. It is evident that increasing the Node count while maintaining a Reference length under approximately 1000 tends to degrade performance, whereas increasing the Node count with a Reference length over 1000 enhances performance. This suggests that longer References complicate model comprehension, but additional structured expression aids understanding when Reference lengths are extensive. Conversely, when Reference lengths are shorter, introducing more structured expressions may introduce noise and reduce comprehension.
Fig. 5 displays the impact of different Reference lengths and Node counts on GPT4-Turbo across various languages and tasks. Significant impacts are noted in Node.Depth, PathCompose, and Statistics tasks, as well as in LaTeX, Markdown, and Tree languages, where increasing Reference length and Node count adversely affects performance. This could be due to the absence of complex structural data in these real-world applications, suggesting that training data did not contain excessive complexity, making additional nodes and References increase the complexity LLMs must manage, thereby affecting accuracy and efficiency. However, in the PathWalk task, while increasing the Node count degrades performance, a greater number of References exacerbates this effect by causing information overload and complicating path tracking. Conversely, increasing Reference length and Node count showed minimal impact on JSON, XML, and YAML, which is attributed to the frequent use and, thus, better generalization of GPT4-Turbo in these well-trained languages.
5.2 Performance Analysis of Finetuned Large Language Models
This section analyzes the enhancement of LLMs’ understanding of structure-rich text through finetuning. Fig. 7 and Fig. 7 illustrate the performance changes under different finetuning strategies. Notably, the initial accuracy of the Vanilla Llama-7B-v1 is 0%, which is reflected by the y-axis starting at 0% at 0 steps in all figures.
5.2.1 Finetuning Enhance LLM’s Performance in All Languages
As depicted in Fig. 7, finetuning improves the models’ comprehension of all structured languages. Although the base finetuning strategy may compromise some capabilities in general free-text language, it notably enhances the understanding of structured data, particularly in YAML and Tree languages, achieving performance comparable to GPT4-Turbo. This indirectly validates the rationality of the task settings for Tree language, as appropriate training enables LLMs to master this custom-designed language effectively. In contrast, while preserving capabilities in free-text language, the mix-fine-tuning strategy offers limited enhancement in the comprehension of structured data, with minimal improvements in Tree language. This may be due to the significant differences between the representation of Tree language and general text, making it challenging to enhance capabilities in both areas simultaneously with a 7B parameter model.
5.2.2 Not All Tasks can be Enhanced by FineTuning
Fig. 7 shows that changes in task performance are more pronounced than changes in language performance. Base finetuning provides the most substantial and consistent performance improvements across most tasks. As training progresses, model performance generally shows a continuous improvement trend. However, other finetuning methods do not correlate positively with increased training steps and performance enhancements in certain tasks, such as Join and Statistics. Additionally, tasks like Syntax quickly stabilize after a few training rounds, indicating that LLMs can rapidly adapt to and understand the grammatical rules of specific structured languages. This swift convergence may highlight the direct impact of finetuning strategies on the models’ comprehension of specific language structures.
5.3 Case Study
We illustrate the assessment setup and specific evaluation methods of StrucText-Eval through two exemplary cases in Fig. 8.
In the Text Retrieval task under the JSON language, the model is required to extract the most deeply nested objects from a complex nested structure and directly output the dictionary type in free-text format. The GPT4-Turbo model accurately identified the objects that met the criteria and matched the standard answer precisely, demonstrating its profound understanding of structured text. Although the Minimax model’s response was also judged as correct, it did not fully adhere to the preset output requirements. Given that many studies have explored the issue of compliance with output formats , our study focuses on assessing the LLM’s understanding of structured text, thereby considering Minimax’s response acceptable.
In the Join task under SQL language, the model is required to merge two tables and perform joint reasoning to output the number of records that meet specific criteria. The reference data was presented in a format that facilitates visualization. The GPT4-Turbo model, which without any further finetuning, failed to answer this question correctly. However, the model that underwent Base Finetuning exhibited progressive improvement: from initially incorrect responses at 300 training steps to gradually understanding the output format requirements at 1200 steps, and finally accurately solving the problem at 5100 steps. This progression illustrates the model’s incremental enhancement in comprehending complex SQL queries and database structures through appropriate finetuning.
6 Conclusion
We believe that the capability to directly interpret structural-rich text in a free-text format is an essential skill all LLMs require. In response, we have developed StrucText-Eval to evaluate this capability of LLMs. Our findings indicate that the proficiency of current LLMs in training on these structural-rich texts varies depending on user frequency, leading to markedly different outcomes when the same tasks are performed in various languages. Currently, LLMs’ understanding of structural-rich texts remains superficially tied to the training data, and these models lack a profound understanding of the structure itself. This deficiency becomes evident when LLMs encounter complex structures composed of common languages or need to parse structural-rich text by custom languages, resulting in significant performance degradation. Direct training on the dataset provided by StrucText-Eval effectively enhances the LLM’s understanding of structural-rich text. However, training solely on the StrucText-Eval dataset does not allow achieving GPT4-Turbo’s comprehensive capabilities on the test set, necessitating the collection of more extensive data and the optimization of training methodologies.
7 Limitation
This paper focuses on evaluating and enhancing the comprehension LLM’s abilities on structure-rich text by designing a benchmark called StrucText-Eval. However, StrucText-Eval includes only eight types of structured languages and encompasses a total of 29 different tasks. Given the vast array of actual structured languages and the myriad methodologies employed beyond these 29 types, StrucText-Eval can only partially represent the LLMs’ capacity to understand structure-rich text. Additionally, due to regional restrictions, we are unable to utilize some highly effective baseline LLMs, such as Gemini and Claude. Therefore, the conclusions drawn in this paper are based on the assumption that GPT-4 and GPT-4 Turbo represent the top-tier LLMs now.
8 Ethical Concern
We contend that this article is devoid of ethical concerns for several reasons:
1.
Nature of StrucText-Eval Content: StrucText-Eval is primarily composed of structured language syntax and some nonsensical strings, which do not present potential ethical issues such as gender bias or racial discrimination.
2.
Objective Presentation of Experimental Results: The experimental results pertaining to StrucText-Eval objectively demonstrate the comprehension abilities of various large models on structure-rich text included in the benchmark. We have thoroughly validated the outputs and assessment details of the models to ensure that the entire evaluation adheres to the experimental setup and maintains objectivity.
3.
Completion of Manual Tasks: All manual tasks associated with this study were conducted by the authors themselves, thereby eliminating any issues of unfair labor practices or unethical cost imposition.
Appendix A Detail about Manual Works
This paper involves the manual works in writing Question Templates and the acquisition of human performance on StrucText-Eval-Hard. All manual works are carried out by the authors of this paper, so there is no payment for the works. The three authors collectively completed the writing and verification of 29 Question Templates, and all templates along with the dataset have been made publicly available on the same website. Moreover, each of the three authors provided responses to 500 identical questions in StrucText-Eval-Hard, with each author dedicating approximately 17 hours. Thus, the human performance results presented in Table 2 are calculated based on the average scores across these 1500 responses.
Appendix B Statistics to StrucText-Eval
The detailed statistic information about StrucText-Eval is illustrated in Tab. 3.
Category | Task | Difficulty | #Samples | #Input | #Instruction | #Answer | #Node |
JSON | PathCompose | hard | 104 | 787.1 | 101.4 | 32.4 | 3 |
simple | 104 | 116.3 | 101.0 | 17.6 | 2 | ||
PathWalk | hard | 104 | 787.1 | 38.0 | 1.1 | 3 | |
simple | 104 | 116.3 | 38.0 | 1.0 | 2 | ||
Syntax | hard | 104 | 736.0 | 148.0 | 4.0 | 3 | |
simple | 104 | 109.0 | 148.0 | 4.0 | 2 | ||
TextRetrieval | hard | 124 | 786.6 | 105.1 | 156.4 | 3 | |
simple | 104 | 116.3 | 94.1 | 75.3 | 2 | ||
TextRetrieval_1∗ | hard | 104 | 787.1 | 161.0 | 315.4 | 3 | |
simple | 104 | 116.3 | 161.0 | 32.2 | 2 | ||
LaTeX | PathWalk | hard | 104 | 2726.1 | 179.1 | 484.0 | 2 |
simple | 104 | 277.2 | 160.3 | 145.1 | 1 | ||
TextRetrieval | hard | 104 | 2726.1 | 74.0 | 54.2 | 2 | |
simple | 104 | 277.2 | 74.0 | 16.1 | 1 | ||
TextRetrieval_1 | hard | 104 | 2726.1 | 62.0 | 57.1 | 2 | |
simple | 104 | 277.2 | 62.0 | 13.3 | 1 | ||
Markdown | PathWalk | hard | 104 | 2355.7 | 184.8 | 370.9 | 3 |
simple | 104 | 228.5 | 162.5 | 114.5 | 2 | ||
TextRetrieval | hard | 104 | 2355.7 | 73.0 | 52.5 | 3 | |
simple | 104 | 228.5 | 73.0 | 12.4 | 2 | ||
TextRetrieval_1 | hard | 104 | 2355.7 | 62.0 | 52.1 | 3 | |
simple | 104 | 228.5 | 62.0 | 12.2 | 2 | ||
ORG | PathWalk | hard | 104 | 2252.7 | 178.6 | 373.3 | 3 |
simple | 104 | 200.1 | 156.5 | 104.5 | 2 | ||
TextRetrieval | hard | 104 | 2252.7 | 73.0 | 52.0 | 3 | |
simple | 104 | 200.1 | 73.0 | 15.2 | 2 | ||
TextRetrieval_1 | hard | 104 | 2252.7 | 62.0 | 54.0 | 3 | |
simple | 104 | 200.1 | 62.0 | 12.3 | 2 | ||
SQL | SQL.Join | hard | 104 | 798.6 | 51.0 | 1.0 | 0 |
simple | 104 | 488.3 | 51.0 | 1.0 | 0 | ||
Statistics | hard | 104 | 798.6 | 49.9 | 1.3 | 0 | |
simple | 104 | 488.3 | 49.9 | 1.0 | 0 | ||
Statistics_1 | hard | 104 | 798.6 | 27.0 | 1.1 | 0 | |
simple | 104 | 488.3 | 27.0 | 1.0 | 0 | ||
TextRetrieval | hard | 104 | 798.6 | 43.8 | 4.3 | 0 | |
simple | 104 | 488.3 | 43.6 | 3.9 | 0 | ||
Tree | PathCompose | hard | 104 | 1745.7 | 84.0 | 25.0 | 8 |
simple | 104 | 555.6 | 83.8 | 11.3 | 4 | ||
Tree.Depth | hard | 104 | 1745.7 | 68.0 | 1.0 | 8 | |
simple | 104 | 555.6 | 67.9 | 1.0 | 4 | ||
Tree.Height | hard | 104 | 1745.7 | 156.0 | 1.0 | 8 | |
simple | 104 | 555.6 | 156.0 | 1.0 | 4 | ||
XML | Syntax | hard | 104 | 2135.7 | 147.0 | 4.0 | 5 |
simple | 104 | 2338.0 | 147.0 | 4.0 | 3 | ||
TextRetrieval | hard | 104 | 2621.6 | 54.5 | 17.7 | 5 | |
simple | 104 | 2910.4 | 52.5 | 11.9 | 3 | ||
TextRetrieval_1 | hard | 104 | 2621.6 | 81.9 | 110.4 | 5 | |
simple | 104 | 2910.4 | 83.8 | 132.9 | 3 | ||
YAML | PathCompose | hard | 104 | 631.7 | 101.4 | 32.4 | 6 |
simple | 104 | 82.2 | 101.0 | 19.6 | 4 | ||
PathWalk | hard | 104 | 631.7 | 38.0 | 1.0 | 6 | |
simple | 104 | 82.2 | 38.0 | 1.0 | 4 | ||
Syntax | hard | 104 | 725.8 | 148.0 | 4.0 | 6 | |
simple | 104 | 94.6 | 148.0 | 4.0 | 4 | ||
TextRetrieval | hard | 104 | 631.7 | 94.3 | 111.9 | 6 | |
simple | 104 | 82.2 | 94.0 | 51.9 | 4 | ||
TextRetrieval_1 | hard | 104 | 631.7 | 161.0 | 224.2 | 6 | |
simple | 104 | 82.2 | 161.0 | 22.1 | 4 |
Appendix C Detail Setting
All experiments and training process is carried out on a three 3090 GPUs service. The setting of API calling and finetuning is illustrated in Tab. 4 And the training loss is illustrated in Fig. 9(d).
Random Seed | ||||
torch.manual_seed | torch.cuda.manual_seed_all | numpy.random.seed | random.seed | torch.backends.cudnn.deterministirc |
42 | 42 | 42 | 42 | True |
AutoCausalLM | ||||
temperature | top_p | top_k | num_beams | max_new_token |
0.95 | 0.95 | 5 | 2 | 1 |
Training Detail | ||||
Batch Size | Max Length | All steps | Warm_up Steps | Learning Rate |
64 | 1024 | 5800 | 100 | 1e-5 |
Appendix D Detail Prompt
The prompts used in the experiment can be categorized into three types: Example of Base Prompts are shown in Tab. 5. Example of CoT Prompts are shown in Tab. 6. Example of Few-Shot Prompts are shown in Tab. 7. Example of Rule Hints are shown in Tab. 8.
# -*- coding: utf-8 -*- |
Variables: |
!<INPUT 0>! – Language |
!<INPUT 1>! – Question |
!<INPUT 2>! – Reference |
!<INPUT 3>! – Requirement |
<commentblockmarker>###</commentblockmarker> |
you are a !<INPUT 0>! file parser, you are required to answer questions pertaining to the given !<INPUT 0>! file. |
### Question: |
!<INPUT 1>! |
### Reference: |
!<INPUT 2>! |
### Requirement: |
!<INPUT 3>! |
Please follow the format below for your output: |
### Answer: |
xxxxx |
# -*- coding: utf-8 -*- |
Variables: |
!<INPUT 0>! – Language |
!<INPUT 1>! – Question |
!<INPUT 2>! – Reference |
!<INPUT 3>! – Requirement |
<commentblockmarker>###</commentblockmarker> |
you are a !<INPUT 0>! file parser, you are required to answer questions pertaining to the given !<INPUT 0>! file. |
### Question: |
!<INPUT 1>! |
### Reference: |
!<INPUT 2>! |
### Requirement: |
!<INPUT 3>! |
Please follow the format below for your output: |
### Reasoning Prcess: |
xxxx |
### Answer: |
xxxxx |
# -*- coding: utf-8 -*- |
Variables: |
!<INPUT 0>! – Language |
!<INPUT 1>! – Demonstration |
!<INPUT 2>! – Question |
!<INPUT 3>! – Reference |
!<INPUT 4>! – Requirement |
<commentblockmarker>###</commentblockmarker> |
you are a !<INPUT 0>! file parser, you are required to answer questions pertaining to the given !<INPUT 0>! file. |
### Demonstration: |
!<INPUT 1>! |
### Question: |
!<INPUT 2>! |
### Reference: |
!<INPUT 3>! |
### Requirement: |
!<INPUT 4>! |
Please follow the format below for your output: |
### Answer: |
xxxxx |
# -*- coding: utf-8 -*- |
Variables: |
!<INPUT 0>! – Language |
!<INPUT 1>! – Question |
!<INPUT 2>! – Reference |
!<INPUT 3>! – Requirement |
!<INPUT 4>! – Rule Hint |
<commentblockmarker>###</commentblockmarker> |
you are a !<INPUT 0>! file parser, you are required to answer questions pertaining to the given !<INPUT 0>! file. |
### Question: |
!<INPUT 1>! |
### Reference: |
!<INPUT 2>! |
### Requirement: |
!<INPUT 3>! |
### Rule Hint: |
!<INPUT 4>! |
Please follow the format below for your output: |
### Answer: |
xxxxx |
Appendix E Examples for All Languages & Tasks
In this section, we provide detailed examples for each language we discuss, illustrating how specific tasks are executed within those languages. These examples are meant to offer clear insights into the application and utility of each language in various contexts. Through these demonstrations, readers can better understand the unique features and capabilities of each language when applied to different tasks.
E.1 Tree
See Figure 10.
E.2 Tabular
See Figure 11.
E.3 JSON
See Figure 12.
E.4 YAML
See Figure 13.
E.5 XML
See Figure 14.
E.6 LaTeX
See Figure 15.
E.7 Markdown
See Figure 16.
E.8 Org
See Figure 17.
Appendix F Rules & Rule Hints
We list all the rules in Regular Express in this section, and list all the hints for these rules in Lis. 1.
F.1 Tree
We build tree-structured input as a list of edges in a tree, in a format of “father->child”, separated by newline.
F.2 Tabular
Formally, input texts are classified as tabular data given that they are composed of a list of newline separated lines, each of which is a list of text cells delimited by comma.
F.3 JSON
Due to the inherit hierarchy structure of Object Notations, we adopted a recursive scheme to define our input texts.
F.4 YAML
The rules for constructing YAML and XML input are similarly recursive.
F.5 XML
F.6 LaTeX
In LaTeX input texts, we include textbf and includegraphics commands to accommodate for the text retrieval tasks. The headings serve as anchors for structure traversal.
F.7 Markdown
In markdown input texts, the syntax counterparts for heading, text face and including figure are employed in our dataset.
F.8 Org
In Org input texts, the syntax is obtained from JSON construction rules by replacing the markups for heading, including figures and bold font face.
Appendix G More Evaluation Result
The RougeL Score is illustrated in Tab. G.
Model | Prompt | Languages | Tasks | all | ||||||||||||||
[…]
Ernie-Turbo | Base | 1.30 | 1.40 | 1.00 | 0.90 | 1.00 | 0.50 | 4.00 | 7.00 | 0.10 | 0.60 | 0.00 | 4.70 | 0.20 | 1.20 | 0.70 | 0.70 | 2.40 |
Ernie-Turbo | CoT | 1.10 | 1.20 | 0.90 | 0.90 | 1.20 | 0.90 | 2.00 | 5.70 | 0.10 | 1.10 | 0.00 | 7.20 | 1.50 | 3.10 | 3.00 | 2.30 | 2.00 |
Ernie-Turbo | 1-shot | 11.00 | 16.90 | 19.60 | 17.00 | 3.10 | 1.40 | 1.30 | 12.20 | 0.80 | 13.80 | 0.00 | 16.90 | 1.20 | 0.90 | 2.00 | 0.70 | 10.20 |
Ernie-Turbo | 2-shot | 7.00 | 24.60 | 21.70 | 17.50 | 3.10 | 1.60 | 8.60 | 7.50 | 0.40 | 16.80 | 0.00 | 16.50 | 2.90 | 0.20 | 2.20 | 1.70 | 10.60 |
Ernie-Turbo | Rule Hint | 2.50 | 6.60 | 8.00 | 9.20 | 0.90 | 0.70 | 1.40 | 4.00 | 0.00 | 7.80 | 0.00 | 5.50 | 0.20 | 0.90 | 1.30 | 0.60 | 3.90 |
[…]
Minimax | CoT | 15.50 | 37.90 | 33.80 | 29.90 | 2.30 | 2.20 | 5.70 | 18.20 | 6.20 | 47.30 | 8.40 | 46.60 | 3.60 | 8.30 | 6.30 | 0.00 | 6.60 |
Minimax | 1-shot | 19.50 | 33.30 | 33.10 | 32.80 | 6.30 | 4.30 | 14.10 | 23.90 | 3.10 | 30.70 | 4.40 | 30.40 | 3.30 | 4.20 | 4.70 | 6.90 | 20.50 |
Minimax | 2-shot | 23.60 | 38.50 | 34.80 | 35.50 | 8.60 | 3.70 | 13.30 | 25.30 | 4.40 | 32.80 | 3.40 | 34.00 | 6.70 | 6.10 | 1.60 | 6.90 | 22.60 |
Minimax | Rule Hint | 12.90 | 18.10 | 19.80 | 20.80 | 2.20 | 0.60 | 5.10 | 11.70 | 0.00 | 16.70 | 3.60 | 17.40 | 0.40 | 1.30 | 1.90 | 0.00 | 11.20 |
[…]
Spark | CoT | 0.40 | 0.20 | 6.10 | 0.20 | 0.20 | 0.00 | 0.10 | 0.50 | 0.00 | 6.00 | 0.20 | 1.50 | 0.60 | 0.20 | 0.10 | 0.10 | 0.90 |
Spark | 1-shot | 4.10 | 2.00 | 1.40 | 8.50 | 0.10 | 0.10 | 0.30 | 2.00 | 0.00 | 4.90 | 0.50 | 3.20 | 0.00 | 0.00 | 0.20 | 0.20 | 2.30 |
Spark | 2-shot | 1.00 | 0.50 | 0.50 | 0.60 | 0.40 | 0.10 | 0.20 | 1.10 | 0.00 | 0.60 | 0.10 | 1.00 | 0.00 | 0.10 | 0.10 | 0.10 | 0.60 |
Spark | Rule Hint | 0.40 | 1.20 | 1.00 | 4.00 | 0.30 | 0.10 | 0.70 | 1.30 | 0.00 | 2.60 | 0.20 | 1.20 | 0.30 | 0.30 | 0.10 | 0.10 | 1.00 |
[…]
GPT4 | CoT | 18.40 | 39.90 | 38.50 | 39.60 | 7.20 | 4.40 | 8.60 | 20.70 | 6.20 | 30.80 | 2.10 | 31.50 | 0.90 | 9.90 | 4.70 | 2.00 | 21.20 |
GPT4 | 1-shot | 22.80 | 38.50 | 35.50 | 38.90 | 11.30 | 5.50 | 18.50 | 26.00 | 9.40 | 34.30 | 4.00 | 34.50 | 10.00 | 9.80 | 4.70 | 3.40 | 23.90 |
GPT4 | 2-shot | 23.10 | 38.70 | 38.20 | 36.70 | 14.40 | 5.40 | 17.50 | 26.30 | 8.40 | 35.10 | 3.40 | 34.60 | 13.30 | 14.50 | 6.20 | 4.30 | 24.60 |
GPT4 | Rule Hint | 13.50 | 22.20 | 21.20 | 23.80 | 3.20 | 1.80 | 5.00 | 13.70 | 2.70 | 19.10 | 2.20 | 19.20 | 1.60 | 2.70 | 2.90 | 0.40 | 12.70 |
[…]
GPT4-Turbo | CoT | 8.70 | 31.40 | 30.00 | 27.30 | 2.00 | 1.50 | 7.80 | 13.80 | 2.30 | 45.40 | 3.10 | 43.60 | 2.30 | 3.80 | 4.20 | 1.60 | 13.90 |
GPT4-Turbo | 1-shot | 21.30 | 35.90 | 35.50 | 36.70 | 8.70 | 5.90 | 17.90 | 25.00 | 7.50 | 33.20 | 1.50 | 33.70 | 1.50 | 9.40 | 8.20 | 2.50 | 22.80 |
GPT4-Turbo | 2-shot | 21.70 | 37.50 | 36.30 | 36.70 | 11.10 | 5.10 | 16.90 | 25.30 | 8.50 | 33.40 | 1.50 | 34.10 | 1.90 | 14.00 | 5.00 | 2.10 | 23.30 |
GPT4-Turbo | Rule Hint | 9.70 | 20.70 | 21.90 | 21.90 | 2.00 | 0.60 | 3.70 | 12.20 | 0.70 | 16.70 | 1.60 | 17.40 | 1.50 | 2.00 | 1.20 | 0.20 | 11.10 |